Some bad and some good news for this report!
BSM_Berry is operational, but with a loaner camera from Greg Bolt. The original ST-8XME camera suffered a failure and has been returned to SBIG for repairs.
BSM_Hamren is off-line. Its ST8-XME camera has suffered a catastrophic failure. Dick Post has offered to donate an ST8-XME camera for replacement, which we hope to ship to Mike Linnolt in a few weeks.
BSM_NH is operational with a new Win10 computer (kindly configured by Ken Menzies), as well as a new camera (the ZWO ASI-183mm with an 8-position filter wheel). We're testing this camera as a possible replacement for the aging SBIG cameras. It offers a number of improvements, including higher QE and the ability to go to much brighter stars. There are a couple of software issues remaining in this new configuration, but I expect data to be flowing in a week or two. A new Paramount ME mount, on long-term loan from the Tzec Maun Foundation (thanks to Dick Post for spearheading this donation effort), has arrived and is being tested in the lab.
BSM_NM is operational. We moved it to a new pier at Bill Stein's observatory, and had the camera refurbished over the summer monsoon shutdown.
BSM_S is operational.
SRO was shut down for the monsoon through mid-September. When restarting, a couple of dome problems occurred that are being fixed. Everything seems to be working now, and so new data should be available soon.
OC61 is operational after the mirror realuminization in August and a number of dome/slit problems in September. Ken Menzies is configuring the backup Win10 computer with the latest releases of all software. Sometime soon, we'll switch computers and test.
TMO61 is probably operational, but I haven't heard from the telescope manager lately. The weather has been poor due to an extended monsoon.
Mike Nicholas will be presenting a poster on the Bright Star Monitor network at the Fall meeting.